# Performance of a Superconducting Quantum Battery

### v1

@@ Abstract
Finding a quantum battery model that demonstrates a quantum advantage while remaining feasible for experimental production is a considerable challenge. In this paper, we introduce a superconducting quantum battery (SQB) model that exhibits such an advantage. The model consists of two coupled superconducting qubits that interact during the unitary charging process while remaining in equilibrium with a thermal reservoir. We first describe the model, provide evidence of the quantum advantage, and then discuss the fabrication process of the battery using superconducting qubits. Importantly, we derive analytical expressions for the ergotropy, instantaneous power, and capacity of the SQB, as well as their connection to quantum coherence. We demonstrate that leveraging the collective effects of Josephson energies and the coupling energy between qubits allows for optimization, resulting in improved energy redistribution and a significant enhancement in charging efficiency. This work highlights the complexities of tuning system parameters, which increase the potential for work extraction from the quantum battery, thereby providing a deeper understanding of the charging mechanisms involved. These findings can be applied to superconducting quantum circuit battery architectures, underscoring the feasibility of efficient energy storage in these systems. Our results pave the way for proposals of new superconducting devices capable of storing extractable work, emphasizing their potential for efficient energy storage.
@@ I Introduction
Over the last twenty years, we have witnessed significant progress in quantum technologies, ranging from fundamental physics to commercial applications. This advancement has led to important developments in energy transfer mechanisms within the quantum realm and the role of quantum resources in microscopic devices [1, 2, 3, 4, 5]. By focusing on these advancements, one can see various physical implementations emerging as leading technologies, such as quantum batteries [6, 7, 8, 9] and superconducting circuits [10, 11, 12]. While it is still too early to determine which of these will form the foundation of tomorrow’s quantum batteries [6, 7, 8, 9], superconducting qubits are arguably at the forefront [13]. Research in this area has progressed beyond academia and is now attracting significant interest from major technology companies, including IBM, underscoring its impact on the field.
@@
Optimizing charging processes is a core objective in quantum battery research. Effective quantum control techniques are essential for enabling faster and more reliable charging. Recently, many researchers studying the charging of quantum batteries have shown that quantum coherence and entanglement significantly enhance charging speed and energy storage efficiency [25, 26]. Furthermore, research on the scalability of quantum battery performance with system size reveals that collective quantum effects enable enhanced energy extraction beyond classical capabilities [27, 28, 29, 30, 31, 32, 33]. However, the charging protocols often rely on specific initial battery states. This dependence on the state of the quantum battery remains relatively unexplored. In this work, along these lines, we address and highlight a superconducting quantum battery (SQB) initialized in a thermal state. This battery is constructed from superconducting charge qubits with symmetric Josephson junction energies. These qubits are capacitively coupled and shunted with a large capacitance to suppress charge noise. Known as Transmon qubits, these devices function as artificial atoms, allowing us to harness their Josephson energies—particularly the mutual coupling energy between qubits—to drive ergotropy production. Experimentally, SQBs represent a cutting-edge paradigm in quantum energy storage [13], aiming to exploit quantum correlations and superposition to achieve charging capabilities that surpass those of conventional electrochemical batteries in efficiency and performance.
@@ I.1 Contribution and Motivation
The quest for efficient quantum battery models that exhibit quantum advantages is a pressing challenge in advancing quantum technology. Superconducting qubits, with their established role in quantum computing, offer a promising platform for realizing such quantum batteries. However, developing a feasible design that is both experimentally implementable and optimizes energy transfer requires careful consideration of coherence and interaction dynamics within the system. In this work, we propose a SQB model based on two coupled superconducting qubits. This model demonstrates a significant quantum advantage in energy storage and extraction while remaining achievable for experimental realization. By analyzing the SQB’s ergotropy, power, and capacity, we identify key factors—such as coherence and qubit coupling—that maximize work extraction, illustrating how collective effects from Josephson energies and inter-qubit coupling enhance charging efficiency. This SQB model connects the quantum properties of the system with practical considerations for device fabrication, laying a foundation for developing high-performance energy storage in superconducting quantum circuits.
@@ I.2 Organization
In this work, we focus on the role of coherence in ergotropic work extraction. We begin by briefly reviewing the relevant model of two superconducting charge qubits (SCQs) in Sec. II. The remainder of this section discusses the fabrication of the SQB within this framework and presents the details of the charging process. It also includes fundamental performance indicators for SQBs, such as ergotropy, power, capacity, and coherence. In Sec. III, we qualitatively describe the results and insights gained regarding the performance of the SQB. We summarize our findings and discuss potential future research directions in the final section, Sec. IV.
@@ II Model and performance of SQB
where:
  • $\xi_{c1}$ and $\xi_{c2}$ are the charging energies.
  • $\xi_{1}$ and $\xi_{2}$ are the Josephson energies.
  • $\xi_{c}$ is the mutual coupling energy between the two qubits.
  • $\sigma_{x1}=\sigma_{x}\otimes I$, $\sigma_{x2}=I\otimes\sigma_{x}$, and $\sigma_{zz}=\sigma_{z}\otimes\sigma_{z}$, with $\sigma_{x,z}$ being Pauli matrices and $I$ is the identity matrix.
  • $n_{gj}=\frac{C_{gj}V_{gj}}{2e}$ is the normalized qubit gate charge, where $C_{gj}$ and $V_{gj}$ are the control gate capacitance and voltage, respectively.
For simplicity, calculations are restricted to the degeneracy point, where $n_{g1}=n_{g2}=0.5$, a condition that ensures insensitivity to noise [35]. Under this condition, the model Hamiltonian simplifies to
@@
At $t=0$, if the system is fully discharged, its initial state can be the ground state $\mathcal{R}(0)=|00\rangle\langle 00|$ at absolute zero or Gibbs state at any finite temperature. The effectiveness of a quantum battery depends on a robust charging process; however, these processes are often sensitive to the initial state of the battery. We assume the Gibbs state as the initial state of QB instead of ground state as it is practical experimentally to deal with finite temperature QB than QB at ground state. Therefore, for a system at finite temperature $T$, we introduce the initial state described by the Gibbs thermal state:
@@
Figure 1: Schematic diagram of the charging process of the quantum battery.
@@ II.1 Charging closed SQB with Pauli-X gate
Let us consider a closed SQB, where charging protocols are necessarily unitary. First, we note that for almost density matrix, there exists a unitary map that increases its energy. This can be understood by recognizing that the problem of charging with respect to $\mathcal{H_{B}}$ is equivalent to discharging with respect to $-\mathcal{H_{B}}$. Therefore, the only states that cannot be charged using a unitary protocol are those whose ergotropy [38] with respect to $-\mathcal{H_{B}}$ is zero. From Refs. [39, 40, 41], we conclude that the only states that cannot be charged by any unitary operation are those that are diagonal in the eigenbasis of $\mathcal{H_{B}}$ with eigenvalues that are non-decreasing in energy.
The SQB can be charged by applying a local field pulse in the $x$-direction with pulse strength $\Omega=\Omega(t)$. The expression for the charging Hamiltonian ($\mathcal{H}_{c_{x}}$) in $x$-direction is
@@ II.2 Ergotropy, power, capacity, and quantum coherence
where $\hat{\rho}_{\downarrow}=|0^{\otimes N}\rangle\langle 0^{\otimes N}|$ and $\hat{\rho}_{\uparrow}=|1^{\otimes N}\rangle\langle 1^{\otimes N}|$ represent the ground and excited states of the $N$-partite quantum battery, respectively. For our specific case of two qubits, $N=2$. This measure directly quantifies the energy gap between the quantum battery’s maximum and minimum energy states. Its calculation is straightforward and does not involve time-dependent optimization, making it suitable for both open and closed quantum batterys. As such, $\mathcal{K}$ provides a practical and efficient metric for evaluating quantum battery performance.
@@ III Results and discussion
In Fig. 2, we present the variations of key quantities associated with our model: ergotropy $\mathcal{E}$, instantaneous power $\mathcal{P}$, capacity $\mathcal{K}$, and the $l_{1}$-norm of coherence $\mathcal{C}_{l_{1}}$ as functions of $\tau=\Omega t$ for different values of $\xi_{2}$. The other parameters are fixed with $\xi_{1}=1.5$, $T=0.5$, and $\xi_{c}=0.05$.
@@
Figure 2(a) depicts the plots of ergotropy, $\mathcal{E}$, as a function of $\tau$. Ergotropy, representing the extractable work from the quantum system, displays oscillatory behavior, signifying energy fluctuations within the system’s quantum state. The amplitude of these oscillations increases with $\xi_{2}$, the Josephson energy of the second qubit. This suggests that larger values of $\xi_{2}$ enhance the system’s ability to store and extract work, making it more efficient for quantum energy storage or harvesting applications. These results imply that tuning the Josephson energy of the second qubit could optimize the quantum battery’s performance in terms of maximizing its stored work.
@@
Figure 2(b) illustrates the instantaneous power, $\mathcal{P}$, which reflects the rate at which energy is either being extracted from or injected into the system. Positive values of $\mathcal{P}$ correspond to charging the quantum battery, while negative values indicate discharging. The oscillatory nature of $\mathcal{P}$ corresponds to the behavior of ergotropy in Fig. 2(a). For instance, during the interval $\tau\in(0;\pi/2)$, $\mathcal{P}>0$, meaning the system is charging the quantum battery. This aligns with the increasing ergotropy over the same interval, which achieves the maximum value for $\tau=\pi/2$. Conversely, for $\tau\in(\pi/2;\pi)$, $\mathcal{P}$ becomes negative, indicating that the system is discharging, as reflected by the decreasing ergotropy. At $\tau=\pi$, ergotropy drops to zero, marking the point where the system has fully discharged, and the sign of $\mathcal{P}$ flips, indicating the start of a new charging cycle. This periodic behavior of charging and discharging continues, governed by the system’s oscillatory dynamics
The dependence of power, $\mathcal{P}$, on $\xi_{2}$ is significant; for larger values of $\xi_{2}$, there is a noticeable increase in the peak values of oscillation. This behavior demonstrates the impact of system parameters on the dynamic performance of energy extraction processes.
Figure 2(c) displays the capacity $\mathcal{K}$ as a function of $\tau$. In contrast to the other quantities, the capacity remains a constant function in the domain of $\tau$. However, the constant value of $\mathcal{K}$ increases with larger $\xi_{2}$, indicating that an enhanced inducative energy allows for greater information transmission capabilities in the system. This behavior suggests a stable performance in terms of capacity, which is advantageous for maintaining effective communication in quantum protocols, regardless of the oscillatory behaviors observed in ergotropy and power.
@@
Figure 2(d) examines the $l_{1}$-norm of coherence $\mathcal{C}_{l_{1}}$. The coherence exhibits a more complex behavior, with clear periodic features reflecting the underlying quantum interference effects. As $\xi_{2}$ increases, the coherence shows enhanced oscillatory behavior, suggesting that increased coupling may facilitate better preservation of quantum coherence in the system. This coherence is essential for maximizing the performance of quantum protocols, further illustrating the importance of parameter selection in the optimization of quantum resources.
@@
Overall, Fig. 2 shows the interplay between ergotropy, instantaneous power, capacity, and coherence. The results indicate that careful tuning of the parameters, particularly $\xi_{2}$, can significantly influence the efficiency of energy extraction and the preservation of quantum coherence, which are critical for the development of advanced quantum technologies.
@@
Figure 3(a) presents the ergotropy $\mathcal{E}$ as a function of $\tau$. While the general oscillatory behavior is similar to that observed in Fig. 2(a), the patterns differ due to the distinct roles played by $\xi_{1}$ and $\xi_{2}$. In Fig. 3(a), $\xi_{1}$, the Josephson energy of the first qubit, is varied while $\xi_{2}$ is held fixed. As $\xi_{1}$ increases, the amplitude of ergotropy oscillations grows, indicating an enhancement in the SQB’s work extraction potential. This result emphasizes the significance of $\xi_{1}$ in determining the system’s energy dynamics, as it directly governs the properties of the first qubit. However, the influence of $\xi_{1}$ is modulated by the coupling energy $\xi_{c}$, which governs the interaction strength between the two qubits. The higher value of $\xi_{c}$ in case of Fig. 3(a) ($\xi_{c}=0.5$) leads to more diverse shapes, reflecting the stronger interplay between the qubits. In contrast, Fig. 2(a) displayed smoother oscillations due to a weaker coupling $\xi_{c}=0.05$.
Figure 3(b) illustrates the instantaneous power $\mathcal{P}$ as a function of $\tau$, showing a similar oscillatory behavior as in Fig. 2(b), but with a notable increase in complexity. The variation of $\xi_{1}$ affects the power dynamics differently than $\xi_{2}$ due to the stronger coupling strength $\xi_{c}$. Here, the non-sinusoidal oscillations become more noticeable, indicating enhanced energy exchange between the two qubits. The peak values of instantaneous power also grow with increasing $\xi_{1}$, reinforcing the role of this parameter in optimizing energy extraction. The higher coupling strength $\xi_{c}$ amplifies the interaction between the qubits, causing their dynamics to be more interdependent and resulting in a more complex power profile compared to Fig. 2(b).
@@
Figure 3(c) presents the capacity $\mathcal{K}$ as a function of $\tau$. As in Fig. 2(c), the capacity is a constant function in the domain of $\tau$, yet its value increases for larger $\xi_{1}$. This stability in capacity, despite oscillations in ergotropy and power, indicates a consistent performance in information transmission capabilities across varying parameter regimes, similar to the observations made in Fig. 2.
Figure 3(d) examines the $l_{1}$-norm of coherence $\mathcal{C}_{l_{1}}$. The coherence showcases oscillatory behavior, paralleling the dynamics observed in Fig. 2(d). As $\xi_{1}$ increases, the oscillations become more significant, indicating that higher values of $\xi_{1}$ enhance the coherence properties of the quantum state. However, this improvement appears to plateau at higher values of $\xi_{1}$, as the corresponding plots converge and nearly overlap, suggesting a saturation effect where further increases in $\xi_{1}$ does not yield to coherence enhancement.
The results in Fig. 3 mirror several trends observed in Fig. 2, demonstrating the key role of parameter selection in optimizing the system’s performance. The coherent interplay between ergotropy, instantaneous power, capacity, and coherence highlights the complex dynamics of energy extraction and information transmission in quantum systems.
In Fig. 4, we explore the variations of ergotropy $\mathcal{E}$, instantaneous power $\mathcal{P}$, capacity $\mathcal{K}$, and the $l_{1}$-norm of coherence $\mathcal{C}_{l_{1}}$ as functions of $\tau=\Omega t$ for varying values of $\xi_{c}$. The remaining parameters are fixed with $\xi_{1}=\xi_{2}=1.5$ and $T=0.1$.

### v2

@@ Abstract
Finding a quantum battery model that demonstrates a quantum advantage while remaining feasible for experimental production is a considerable challenge. Here, a superconducting quantum battery (SQB) model that exhibits such an advantage is introduced. The model consists of two coupled superconducting qubits that interact during the unitary charging process while remaining in equilibrium with a thermal reservoir. First, the model is described, evidence of the quantum advantage is provided, and then the fabrication process of the battery is discussed using superconducting qubits. Analytical expressions for the ergotropy, instantaneous power, and capacity of the SQB, as well as their connection to quantum coherence are derived. It is demonstrated that leveraging the collective effects of Josephson energies and the coupling energy between qubits allows for optimization, resulting in improved energy redistribution and a significant enhancement in charging efficiency. This work highlights the complexities of tuning system parameters, which increase the potential for work extraction from the SQB, providing a deeper understanding of the charging mechanisms involved. These findings can be applied to superconducting quantum circuit battery architectures, underscoring the feasibility of efficient energy storage in these systems. These results pave the way for proposals of new superconducting devices, emphasizing their potential for efficient energy storage.
@@ I Introduction
Over the last twenty years, we have witnessed significant progress in quantum technologies, ranging from fundamental physics to commercial applications. This advancement has led to important developments in energy transfer mechanisms within the quantum realm and the role of quantum resources in microscopic devices [1, 2, 3, 4, 5, 6, 7]. By focusing on these advancements, one can see various physical implementations emerging as leading technologies, such as quantum batteries [8, 9, 10, 11] and superconducting circuits [12, 13, 14]. While it is still too early to determine which of these will form the foundation of tomorrow’s quantum batteries [8, 9, 10, 11], superconducting qubits are arguably at the forefront [15, 16, 17]. Research in this area has progressed beyond academia and is now attracting significant interest from major technology companies, including International Business Machines(IBM) corporation, underscoring its impact on the field.
@@
Optimizing charging processes is a core objective in quantum battery research. Effective quantum control techniques are essential for enabling faster and more reliable charging. Recently, many researchers studying the charging of quantum batteries have shown that quantum coherence and entanglement significantly enhance charging speed and energy storage efficiency [29, 30]. Furthermore, research on the scalability of quantum battery performance with system size reveals that collective quantum effects enable enhanced energy extraction beyond classical capabilities [31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. However, the charging protocols often rely on specific initial battery states. This dependence on the state of the quantum battery remains relatively unexplored. In this work, along these lines, we address and highlight a superconducting quantum battery (SQB) initialized in a thermal state. This battery is constructed from superconducting charge qubits with symmetric Josephson junction energies. These qubits are capacitively coupled and shunted with a large capacitance to suppress charge noise. Known as Transmon qubits, they function as artificial atoms, allowing us to harness their Josephson energies, particularly the mutual coupling energy between qubits, to drive ergotropy production and enhance both energy and ergotropy. SQBs represent a cutting-edge paradigm in quantum energy storage [15], aiming to exploit quantum correlations and superposition to achieve charging capabilities that surpass those of conventional electrochemical batteries in efficiency and performance.
@@ I.1 Contribution and Motivation
The quest for efficient quantum battery models that exhibit quantum advantages is a challenge in advancing quantum technology. Superconducting qubits, with their established role in quantum computing, offer a promising platform for realizing quantum batteries. However, developing a feasible design that optimizes energy transfer requires careful consideration of coherence and interaction dynamics. In this work, we propose an SQB model based on two coupled superconducting qubits. This model demonstrates a significant quantum advantage in energy storage and extraction while remaining experimentally feasible. By analyzing the SQB’s ergotropy, power, and capacity, we identify key factors, such as coherence and qubit coupling, that maximize work extraction, illustrating how collective effects from Josephson energies and inter-qubit coupling enhance charging efficiency. This SQB model connects the quantum properties of the system with practical considerations for device fabrication, laying a foundation for developing high-performance energy storage in superconducting quantum circuits.
Specifically, we focus on a closed SQB, where the charging process is inherently unitary and free from decoherence and dissipation effects. This establishes a theoretical benchmark for the optimal performance of an SQB under ideal coherent dynamics. Although our model includes a thermal reservoir, it does not induce dissipation during the charging process; the reservoir only defines the initial equilibrium state of the qubits, ensuring that the system remains effectively closed throughout the unitary evolution. In practical superconducting architectures, qubits are subject to noise sources such as charge noise, flux noise, and relaxation effects. While these are not explicitly considered in our analysis, they represent an important aspect for future investigations into the robustness of the observed quantum advantage under realistic conditions. This work could be further developed in two main directions. First, we propose a universal method for fast charging of an SQB connected to a driven quantum charger system through controlled pure dephasing of the charger. Second, we present a feasible scheme for an SQB that can implement the proposed cycle and provide typical experimental values for the circuit elements.
In particular, the relationship between decoherence time and charging time is crucial for the efficiency of SQB. If the charging time is significantly shorter than the decoherence time, the system can achieve near-optimal performance before noise-induced energy losses become significant. Conversely, if decoherence effects occur on comparable or shorter timescales, the benefits of quantum coherence and entanglement may diminish. Recent studies on superconducting qubit coherence times suggest that state-of-the-art devices exhibit coherence times ranging from tens to hundreds of microseconds [14], which should be considered when designing practical charging protocols.
Although the current work does not explicitly model decoherence, our results will provide a theoretical benchmark for the optimal performance of an SQB in the absence of noise. Future extensions could incorporate open-system effects using techniques such as Lindblad master equations or stochastic noise modeling to assess the SQB’s robustness against decoherence [41, 42]. Recent advancements in quantum error mitigation and dynamical decoupling techniques [43] offer promising methods to extend coherence times and enhance the viability of SQBs in real-world implementations.
@@ I.2 Organization
In this work, we focus on the role of coherence in ergotropic work extraction. We begin by briefly reviewing the relevant model of two superconducting charge qubits (SCQs) in Sec. II. The remainder of this section discusses the fabrication of the SQB within this framework and presents the details of the charging process. It also includes fundamental performance indicators for SQBs, such as ergotropy, power, capacity, and coherence. In Sec. III, we qualitatively describe the results and insights gained regarding the performance of the SQB. A discussion on the scalability of SQBs is provided in Sec. IV. Experimental feasibility and optimization strategies for SQBs are presented in Sec. V. We summarize our findings and discuss potential future research directions in the final section, Sec. VI.
@@ II Model and performance of SQB
where $\xi_{c1}$ and $\xi_{c2}$ are the charging energies, $\xi_{1}$ and $\xi_{2}$ are the Josephson energies, and $\xi_{c}$ is the mutual coupling energy between the two qubits. Besides, $\sigma_{x(z)1}=\sigma_{x(z)}\otimes I$, $\sigma_{x(z)2}=I\otimes\sigma_{x(z)}$, and $\sigma_{zz}=\sigma_{z}\otimes\sigma_{z}$, with $\sigma_{x,z}$ being Pauli matrices and $I$ is the identity matrix. Also, $n_{gj}=\frac{C_{gj}V_{gj}}{2e}$ is the normalized qubit gate charge, where $C_{gj}$ and $V_{gj}$ are the control gate capacitance and voltage, respectively.
In our analysis, we restrict our calculations to the degeneracy point ($n_{g1}=n_{g2}=1/2$), as this condition minimizes charge noise sensitivity and increases the coherence properties of the superconducting qubits [45]. This choice is a widely adopted strategy in superconducting qubit studies, where operating at the degeneracy point allows for more stable quantum coherence, reducing the impact of environmental fluctuations [48]. By focusing on this regime, we ensure that the system remains robust against low-frequency charge noise, which is one of the dominant decoherence mechanisms in superconducting circuits [49]. However, we acknowledge the importance of investigating the system’s behavior at different operating points, as deviations from the degeneracy point may introduce charge dispersion effects that influence qubit dynamics. Future work will extend this model to analyze how variations in $n_{g1}$ and $n_{g2}$ affect quantum correlations and coherence, providing a more comprehensive understanding of the system’s performance under realistic operating conditions. Such an extension is crucial for assessing the practical feasibility of implementing superconducting qubits in scalable quantum technologies, where external perturbations and fabrication imperfections may lead to operation outside the ideal degeneracy condition [50]. Furthermore, a broader analysis will help characterize the interplay between charge noise and qubit interactions, offering insights into potential mitigation strategies for improving qubit fidelity in experimental implementations.
Under this condition, the Hamiltonian (1) simplifies to
@@
At $t=0$, if the system is fully discharged, its initial state can be the ground state $\mathcal{R}(0)=|00\rangle\langle 00|$ at absolute zero or Gibbs state at any finite temperature. The effectiveness of a quantum battery depends on a robust charging process; however, these processes are often sensitive to the initial state of the battery. We assume the Gibbs state as the initial state of quantum battery instead of ground state as it is practical experimentally to deal with finite temperature quantum battery than quantum battery at ground state. Therefore, for a system at finite temperature $T$, we introduce the initial state described by the Gibbs thermal state:
@@
Figure 1: Schematic circuit of the proposed SQB charging protocol. The system is initially a) prepared in a Gibbs state, consisting of two superconducting qubits at thermal equilibrium. These qubits correspond to the two Josephson junctions and represent the first and second Cooper pair boxes, respectively. Charging of the SQB occurs under a local field pulse $\Omega(t)$, which provides the necessary energy. Panels b) and c) represent the uncharged and charged states of the battery, respectively.
@@ II.1 Charging closed SQB with Pauli-X gate
We consider the charging process of a SQB in a closed quantum system, where charging protocols are necessarily unitary. First, we note that for almost every density matrix, there exists a unitary map that increases its energy. This can be understood by recognizing that the problem of charging with respect to $\mathcal{H_{B}}$ is equivalent to discharging with respect to $-\mathcal{H_{B}}$. Therefore, the only states that cannot be charged using a unitary protocol are those whose ergotropy [51] with respect to $-\mathcal{H_{B}}$ is zero. From Refs. [52, 53, 54], we conclude that the only states that cannot be charged by any unitary operation are those that are diagonal in the eigenbasis of $\mathcal{H_{B}}$ with eigenvalues that are non-decreasing in energy.
The SQB can be charged by applying a local field pulse in the $x$-direction with pulse strength $\Omega=\Omega(t)$. This field transfers energy to charge the SQB. The expression for the charging Hamiltonian $\mathcal{H}_{c_{x}}$ in the $x$-direction is
@@ II.2 Ergotropy, power, capacity, and quantum coherence
where $\hat{\rho}_{\downarrow}=|0^{\otimes N}\rangle\langle 0^{\otimes N}|$ and $\hat{\rho}_{\uparrow}=|1^{\otimes N}\rangle\langle 1^{\otimes N}|$ represent the ground and excited states of the $N$-partite quantum battery, respectively. For our case of two qubits, we have $N=2$. This measure directly quantifies the energy gap between the maximum and minimum energy states of the quantum battery. Its calculation is straightforward and does not require time-dependent optimization, making it suitable for both open and closed quantum batteries. Therefore, $\mathcal{K}$ serves as a practical and efficient metric for evaluating quantum battery performance.
@@ III Results and discussion
In this section, we investigate the charging performance of the SQB under various physical parameter regimes. We compute and plot the ergotropy $\mathcal{E}$, instantaneous power $\mathcal{P}$, capacity $\mathcal{K}$, and the $l_{1}$-norm of coherence $\mathcal{C}_{l_{1}}$. First, we vary the Josephson energies $\xi_{1,2}$ between the qubits, and then we examine the battery performance under different strengths of the mutual coupling energy $\xi_{c}$. Finally, we explore the significance of the mutual coupling energy $\xi_{c}$ in the system and its role in the efficient charging of the SQB. We consider both equal and varying Josephson energies in the SQB model.
In Fig. 2, we present the variations of key quantities associated with our model: ergotropy $\mathcal{E}$, instantaneous power $\mathcal{P}$, capacity $\mathcal{K}$, and the $l_{1}$-norm of coherence $\mathcal{C}_{l_{1}}$ as functions of $\tau=\Omega t$ for different values of $\xi_{2}$. The parameter $\xi_{2}$ represents the Josephson energy of one of the qubits, which influences its individual energy scale and contributes to the overall charging dynamics of the quantum battery. The fixed parameter $\xi_{1}=1.5$ corresponds to the Josephson energy of the second qubit, playing a similar role in defining the system’s energy distribution. The mutual coupling energy $\xi_{c}=0.05$ determines the strength of interaction between the two qubits, impacting both energy transfer efficiency and coherence properties. The temperature parameter $T=0.5$ governs thermal fluctuations, where higher values lead to increased decoherence effects.
To systematically analyze the impact of $\xi_{2}$, we consider four representative values: $\xi_{2}=0.1,0.5,1$, and $2$. These choices span a range from weak ($\xi_{2}=0.1$) to strong ($\xi_{2}=2$) Josephson energy, allowing us to explore different operating regimes of the quantum battery. At lower values of $\xi_{2}$, the energy gap of the qubit is smaller, leading to slower charging dynamics, while at higher values, the system enters a regime of enhanced energy storage and coherence effects. This selection enables us to contrast scenarios where the qubit behaves as a weakly or strongly coupled element within the battery system. A more detailed analysis of these parameter choices is provided in Appendix D.
@@
Figure 2(a) depicts the plots of ergotropy, $\mathcal{E}$, as a function of $\tau$. Ergotropy, representing the extractable work from the quantum system, displays oscillatory behavior, signifying energy fluctuations within the system’s quantum state. The amplitude of these oscillations increases with $\xi_{2}$, the Josephson energy of the second qubit. This suggests that larger values of $\xi_{2}$ enhance the system’s ability to store and extract work, making it more efficient for quantum energy storage. These results imply that tuning the Josephson energy of the second qubit could optimize the quantum battery’s performance in terms of maximizing its stored work.
@@
Figure 2(b) illustrates the instantaneous power, $\mathcal{P}$, which reflects the rate at which energy is either being extracted from or injected into the system. Positive values of $\mathcal{P}$ correspond to charging the quantum battery, while negative values indicate discharging. The oscillatory nature of $\mathcal{P}$ corresponds to the behavior of ergotropy in Fig. 2(a). For instance, during the interval $\tau\in(0;\pi/2)$, $\mathcal{P}>0$, meaning the system is charging the quantum battery. This aligns with the increasing ergotropy over the same interval, which achieves the maximum value for $\tau=\pi/2$. Conversely, for $\tau\in(\pi/2;\pi)$, $\mathcal{P}$ becomes negative, indicating that the system is discharging, as reflected by the decreasing ergotropy. At $\tau=\pi$, ergotropy drops to zero, marking the point where the system has fully discharged, and the sign of $\mathcal{P}$ flips, indicating the start of a new charging cycle. This periodic behavior of charging and discharging continues, governed by the system’s oscillatory dynamics. The dependence of power, $\mathcal{P}$, on $\xi_{2}$ is significant; larger values of $\xi_{2}$ lead to a noticeable increase in peak oscillation values. This behavior demonstrates how system parameters enhance battery charging performance, resulting in a rapid transfer of energy to the battery.
Fig. 2(c) displays the capacity $\mathcal{K}$ as a function of $\tau$. In contrast to the other quantities, the capacity remains a constant function in the domain of $\tau$. However, the constant value of $\mathcal{K}$ increases with larger $\xi_{2}$, indicating that enhanced Josephson energy is more effective in distributing energy uniformly across the battery, leading to a stable increase in the extractable energy. Although most properties of battery capacity can be derived from ergotropy, we argue that battery capacity is a more fundamental quantity, as it remains unchanged during unitary charging or discharging.
@@
Fig. 2(d) examines the $l_{1}$-norm of coherence $\mathcal{C}_{l_{1}}$. The coherence exhibits a more complex behavior, with clear periodic features reflecting the underlying quantum interference effects. As $\xi_{2}$ increases, the coherence shows enhanced oscillatory behavior, suggesting that increased coupling may facilitate better preservation of quantum coherence in the system. This coherence is essential for maximizing the performance of quantum protocols, further illustrating the importance of parameter selection in the optimization of quantum resources.
@@
Overall, Fig. 2 shows the interplay between ergotropy, instantaneous power, capacity, and coherence. The results indicate that careful tuning of the parameters, particularly $\xi_{2}$, can significantly influence both the efficiency of energy extraction and the preservation of quantum coherence, which are critical for achieving high efficiency in quantum batteries.
@@
Fig. 3(a) presents the ergotropy $\mathcal{E}$ as a function of $\tau$. While the general oscillatory behavior is similar to that observed in Fig. 2(a), the patterns differ due to the distinct roles played by $\xi_{1}$ and $\xi_{2}$. In Fig. 3(a), $\xi_{1}$, the Josephson energy of the first qubit, is varied while $\xi_{2}$ is held fixed. As $\xi_{1}$ increases, the amplitude of ergotropy oscillations grows, indicating an enhancement in the SQB’s work extraction potential. This result emphasizes the significance of $\xi_{1}$ in determining the system’s energy dynamics, as it directly governs the properties of the first qubit. However, the influence of $\xi_{1}$ is modulated by the coupling energy $\xi_{c}$, which governs the interaction strength between the two qubits. The higher value of $\xi_{c}$ in Fig. 3(a) ($\xi_{c}=0.5$) results in more diverse shapes, reflecting the stronger interplay between the qubits. In contrast, Fig. 2(a) shows smoother oscillations due to the weaker coupling ($\xi_{c}=0.05$).
Fig. 3(b) illustrates the instantaneous power $\mathcal{P}$ as a function of $\tau$, showing a similar oscillatory behavior as in Fig. 2(b), but with a notable increase in complexity. The variation of $\xi_{1}$ affects the power dynamics differently than $\xi_{2}$ due to the stronger coupling strength $\xi_{c}$. Here, the non-sinusoidal oscillations become more noticeable, indicating enhanced energy exchange between the two qubits. The peak values of instantaneous power also grow with increasing $\xi_{1}$, reinforcing the role of this parameter in optimizing energy extraction. The higher coupling strength $\xi_{c}$ amplifies the interaction between the qubits, causing their dynamics to be more interdependent and resulting in a more complex power profile compared to Fig. 2(b).
@@
Fig. 3(c) shows the capacity $\mathcal{K}$ as a function of $\tau$. As in Fig. 2(c), capacity is a constant function in the range of $\tau$, with higher $\xi_{1}$ generally leading to a greater increase in capacity in the absence of transient oscillations. This stability of capacity, despite fluctuations in ergotropy and power, indicates that an SQB can store and deliver energy in different parameter regimes, as shown by the observations in Fig. 2. However, in our SQB charging model, the system’s capacity is mainly decided by the Josephson energy of the qubit.
Fig. 3(d) examines the $l_{1}$-norm of coherence $\mathcal{C}_{l_{1}}$. The coherence exhibits oscillatory behavior, paralleling the dynamics observed in Fig. 2(d). As $\xi_{1}$ increases, the oscillations become more pronounced, indicating that higher values of $\xi_{1}$ enhance the coherence properties of the quantum state. However, this improvement appears to plateau at higher values of $\xi_{1}$, as the corresponding plots converge and nearly overlap, suggesting a saturation effect where further increases in $\xi_{1}$ do not lead to additional coherence enhancement.
The results in Fig. 3 reflect trends observed in Fig. 2, highlighting the importance of parameter selection for controlling the charging time of the SQB to ensure maximum charging energy. These parameters influence ergotropy, instantaneous power, and coherence oscillator frequency, which are critical for maximizing energy transfer and achieving higher transient capacity values. The interplay among ergotropy, instantaneous power, capacity, and coherence demonstrates that the SQB can store energy more quickly and has a greater maximum extractable work, as quantified by ergotropy.
[…]